OpenStudy (anonymous):

HELP !! Period of y=-20cos(5/3t)+100 ??

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

In radians or degrees?

OpenStudy (amriju):

do u know the period of cos x?

OpenStudy (anonymous):

degree @KeithAfasCalcLover

OpenStudy (anonymous):

I get 216 degrees

OpenStudy (anonymous):

Ill show how

OpenStudy (anonymous):

Find equation of Sine function with amplitude +5 and period is 240

OpenStudy (anonymous):

@KeithAfasCalcLover

OpenStudy (anonymous):

Normally, for a formula of the form \[a*cos\big[k(x-d)\big]+c\] has a period of \[\frac{360^\circ}{k}\] So this one has a k value of 5/3 and therefore the period is: \[\frac{360^\circ}{1}/\frac{5}{3}=\frac{360}{1}*\frac{3}{5}=\frac{1080}{5}=\frac{216}{1}=216\]

OpenStudy (anonymous):

ok thanks can you help with the second one ?

OpenStudy (anonymous):

Haha sure. Sin is similar. Amplitude is the "a" value for the general formula I wrote above. That being said, we also know that the period wich is 360/k needs to be 240 so lets find the k value:

OpenStudy (anonymous):

Ill represent period by the symbol \(\lambda\). \[\lambda=\frac{360}{k} \rightarrow 240=\frac{360}{k} \rightarrow 240k=360 \rightarrow k=\frac{360}{240}=1.5\]

OpenStudy (anonymous):

Therefore the equation is: \(y=5cos[1.5(x-d)]+c\) where \(c\) and \(d\) are whatever you want them to be! Any number for them will still be correct.

OpenStudy (anonymous):

is has to be y=5sin not cos

OpenStudy (anonymous):

@KeithAfasCalcLover ^^

OpenStudy (anonymous):

Oh hahaha yes correct. Sorry

OpenStudy (anonymous):

@KeithAfasCalcLover help with this one

OpenStudy (anonymous):

@genius12 help

OpenStudy (anonymous):

OHH this is fun.

OpenStudy (anonymous):

OK. Give me a sec.

OpenStudy (anonymous):

@cupcakezz i told u for new questions make a new thread. don't post all of them on a single page..

OpenStudy (anonymous):

Ah good point haha post it on a new page, tag me and then ill answer